# Generic properties of minimal surfaces

### v1

@@
{quoting}
[leftmargin=7mm] Abstract  Let $M$ be an open Riemann surface and $n\geq 3$ be an integer. In this paper we establish some generic properties (in Baire category sense) in the space of all conformal minimal immersions $M\to\mathbb{R}^{n}$ endowed with the compact-open topology, pointing out that a generic such immersion is chaotic in many ways. For instance, we show that a generic conformal minimal immersion $u\colon M\to\mathbb{R}^{n}$ is non-proper, almost proper, and $\mathfrak{g}$-complete with respect to any given Riemannian metric $\mathfrak{g}$ in $\mathbb{R}^{n}$. Further, its image $u(M)$ is dense in $\mathbb{R}^{n}$ and disjoint from $\mathbb{Q}^{3}\times\mathbb{R}^{n-3}$, and has infinite area, infinite total curvature, and unbounded curvature on every open set in $\mathbb{R}^{n}$. In case $n=3$, we also prove that a generic conformal minimal immersion $M\to\mathbb{R}^{3}$ has infinite index of stability on every open set in $\mathbb{R}^{3}$.
@@ 1.  Introduction and main results
Throughout the paper, $n\geq 3$ is an integer, $M$ is an open Riemann surface, and ${\rm CMI}(M,\mathbb{R}^{n})$ denotes the space of all conformal minimal immersions $M\to\mathbb{R}^{n}$ endowed with the compact-open topology. A sequence $u_{j}\in{\rm CMI}(M,\mathbb{R}^{n})$, $j\in\mathbb{N}$, converges in this topology to an immersion $u\in{\rm CMI}(M,\mathbb{R}^{n})$ if and only if the sequence of restrictions $u_{j}|_{K}$ to any compact set $K\subset M$ converges uniformly on $K$ to $u|_{K}$ (see, e.g., [7, §VII.2]). In this paper we shall establish several generic properties in ${\rm CMI}(M,\mathbb{R}^{n})$. Some of them are somewhat surprising, and show that a generic conformal minimal immersion $M\to\mathbb{R}^{n}$ has a considerably wild asymptotic behavior. The following are some of the properties we are interested in. We denote by ||$\cdot$|| ⋅ | and $\mathrm{dist}(\cdot,\cdot)$ the standard Euclidean norm and distance in $\mathbb{R}^{n}$, respectively.
@@
[…]
@@
The following result is a consequence of Theorem 1.2(ii) and (viii), and the fact that conformal minimal immersions $M\to\mathbb{R}^{n}$, $n=3,4$, rebuilding every conformal minimal disc are densely non-injective (see Definition 1.1 (d) and Claim 3.1). Recall that an immersion $u\in{\rm CMI}(M,\mathbb{R}^{4})$ has simple double points if for any $p,q\in M$, $p\neq q$, with $u(p)=u(q)$, the tangent planes $du_{p}(T_{p}M)$ and $du_{q}(T_{q}M)$ intersect only at $0\in\mathbb{R}^{4}$, and there are no triple intersections.
@@
    $\mathcal{F}(\cdot,C)\colon{\rm CMI}(M,\mathbb{R}^{n})\to[0,+\infty]$ is continuous for all $C\in E$, where $[0,+\infty]$ has the topology of the extended real line.
@@
In order to establish each of the statements, we shall consider a pair $(E,\mathcal{F})$ as in Claim 2.3 such that $\bigcap_{i\in\mathbb{N}}\Lambda_{i}$ equals the set we are dealing with. Condition (A) in the claim will always be clear, either trivially or from the Cauchy estimates, and we will not discuss it. Thus, the proof reduces to check condition (B). For this, we shall use the following Runge theorem for conformal minimal immersions (see [3, Theorem 5.3] or [5, Theorem 3.6.1]), together with further results in the recently developed theory of approximation for minimal surfaces.
@@
where $\mathrm{dist}_{u}^{\mathfrak{g}}$ denotes the distance on $M$ associated to the pull-back metric $u^{*}\mathfrak{g}$. Let $i\in\mathbb{N}$ and fix $u$, $K$, and $\epsilon$ as in Claim 2.3 (B). Choose $C=bK_{j}\in E$, where $j\in\mathbb{N}$ is so large that $K\subset\mathring{K}_{j}$, and fix a number $\delta>0$ to be specified later. By [2, Lemma 4.1] (see also [5, Lemma 7.3.1]) there is a conformal minimal immersion $\hat{u}:K_{j}\to\mathbb{R}^{n}$ such that $\|\hat{u}-u\|_{K_{j}}<\epsilon/2$ and $\mathrm{dist}_{\hat{u}}(p_{0},C)>\delta$, where $\mathrm{dist}_{\hat{u}}$ denotes the distance function on $M$ associated to the metric $\hat{u}^{*}{ds^{2}}$ induced on $M$ via $\hat{u}$ by the standard Euclidean metric $ds^{2}$ in $\mathbb{R}^{n}$. Since $ds^{2}$ and the given metric $\mathfrak{g}$ are comparable in the compact set $\{x\in\mathbb{R}^{n}\colon\mathrm{dist}(x,u(K_{j}))\leq\epsilon\}\subset\mathbb{R}^{n}$, we can choose $\delta>0$ so large that
@@
[…]
@@
[…]
@@
Let $M$ be an open Riemann surface and $B\subset\mathbb{R}^{3}$ be an open ball. Reasoning as in the proof of Corollary 1.3 (iv), it suffices tho show that the set $X$ of all immersions $u\in{\rm CMI}(M,\mathbb{R}^{3})$ with $u^{-1}(B)\neq\varnothing$ and having infinite index of stability on $u^{-1}(B)$ is a dense $G_{\delta}$ subset. For this, take an exhaustion $K_{j}$, $j\in\mathbb{N}$, of $M$ as in (2.1), and for each $i\in\mathbb{N}$ set
@@
(D) Let $R$ be a compact bordered Riemann surface, and assume that $R$ appears recurrently in $M$, meaning that for any Runge compact set $K\subset M$ there is a smoothly bounded compact domain $L\subset M\setminus K$ being biholomorphic to $R$ such that $K\cup L$ is Runge. Arguing as in the proofs of Theorem 1.2(ii) and Corollary 1.3(ii), one can see that the subspace $\mathrm{CMI}_{R}(M,\mathbb{R}^{n})\subset\mathrm{CMI}(M,\mathbb{R}^{n})$ of immersions rebuilding every conformal minimal immersion $R\to\mathbb{R}^{n}$ is residual (cf. Definition 1.1(c)). Likewise, if $R_{j}$, $j\in\mathbb{N}$, is a sequence of compact bordered Riemann surfaces appearing recurrently in $M$, [^4] By conformal surgery theory, one can construct open Riemann surfaces $M$ such that $R_{j}$ appears recurrently in $M$ for all $j\in\mathbb{N}$. then $\bigcap_{j\in\mathbb{N}}\mathrm{CMI}_{R_{j}}(M,\mathbb{R}^{n})$ is still residual in $\mathrm{CMI}(M,\mathbb{R}^{n})$.
@@ Acknowledgements
This research was partially supported by the State Research Agency (AEI) via the grants no. PID2020-117868GB-I00 and PID2023-150727NB-I00, and the “Maria de Maeztu” Excellence Unit IMAG, reference CEX2020-001105-M, funded by MCIN/AEI/10.13039/501100011033/, Spain.

### v2

@@
  • Abstract  Let $M$ be an open Riemann surface and $n\geq 3$ be an integer. In this paper we establish some generic properties (in Baire category sense) in the space of all conformal minimal immersions $M\to\mathbb{R}^{n}$ endowed with the compact-open topology, pointing out that a generic such immersion is chaotic in many ways. For instance, we show that a generic conformal minimal immersion $u\colon M\to\mathbb{R}^{n}$ is non-proper, almost proper, and $\mathfrak{g}$-complete with respect to any given Riemannian metric $\mathfrak{g}$ in $\mathbb{R}^{n}$. Further, its image $u(M)$ is dense in $\mathbb{R}^{n}$ and disjoint from $\mathbb{Q}^{3}\times\mathbb{R}^{n-3}$, and has infinite area, infinite total curvature, and unbounded curvature on every open set in $\mathbb{R}^{n}$. In case $n=3$, we also prove that a generic conformal minimal immersion $M\to\mathbb{R}^{3}$ has infinite index of stability on every open set in $\mathbb{R}^{3}$.
@@ 1. Introduction and main results
Throughout the paper, $n\geq 3$ is an integer, $M$ is an open Riemann surface, and ${\rm CMI}(M,\mathbb{R}^{n})$ denotes the space of all conformal minimal immersions $M\to\mathbb{R}^{n}$ endowed with the compact-open topology. A sequence $u_{j}\in{\rm CMI}(M,\mathbb{R}^{n})$, $j\in\mathbb{N}$, converges in this topology to an immersion $u\in{\rm CMI}(M,\mathbb{R}^{n})$ if and only if the sequence of restrictions $u_{j}|_{K}$ to any compact set $K\subset M$ converges uniformly on $K$ to $u|_{K}$ (see, e.g., [7, §VII.2]). In this paper we shall establish several generic properties in ${\rm CMI}(M,\mathbb{R}^{n})$. Some of them are somewhat surprising, and show that a generic conformal minimal immersion $M\to\mathbb{R}^{n}$ has a considerably wild asymptotic behavior. The following are some of the properties we are interested in. We denote by $|\cdot|$ and $\mathrm{dist}(\cdot,\cdot)$ the standard Euclidean norm and distance in $\mathbb{R}^{n}$, respectively.
@@
(a) Given a Riemannian metric $\mathfrak{g}$ in $\mathbb{R}^{n}$, an immersion $u\in{\rm CMI}(M,\mathbb{R}^{n})$ is $\mathfrak{g}$-complete if the pull-back metric $u^{*}\mathfrak{g}$ induced on $M$ by $\mathfrak{g}$ via $u$ is complete in the classical sense: $u\circ\gamma$ has infinite $\mathfrak{g}$-length for every divergent path $\gamma:[0,1)\to M$, that is, $\int_{0}^{1}|(u\circ\gamma)^{\prime}(t)|_{\mathfrak{g}}dt=+\infty$. The immersion $u$ is complete if it is $ds^{2}$-complete, where $ds^{2}$ denotes the standard Euclidean metric in $\mathbb{R}^{n}$.
@@
The following result is a consequence of Theorem 1.2(ii) and (viii), and the fact that conformal minimal immersions $M\to\mathbb{R}^{n}$, $n=3,4$, rebuilding every conformal minimal disc are densely non-injective (see Definition 1.1 (d) and Claim 3.1). Recall that an immersion $u\in{\rm CMI}(M,\mathbb{R}^{4})$ has simple double points if for any $\in M, pqp\neq q$, with $u(p)=u(q)$, the tangent planes $du_{p}(T_{p}M)$ and $du_{q}(T_{q}M)$ intersect only at $0\in\mathbb{R}^{4}$, and there are no triple intersections.
@@
    $\mathcal{F}(\cdot,C)\colon{\rm CMI}(M,\mathbb{R}^{n})\to[0,+\infty]$ is continuous for all $C\in E$, where $0,+\infty]$ has the topology of the extended real line.
@@
In order to establish each of the statements, we shall consider a pair $(E,\mathcal{F})$ as in Claim 2.3 such that $\bigcap_{i\in\mathbb{N}}\Lambda_{i}$ equals the set we are dealing with. Condition (A) in the claim will always be clear, either trivially or from a straightforward application of the Cauchy estimates, and we will not discuss it (except when proving statement (i), serving to exemplify how the Cauchy estimates are used). Thus, the proof reduces to check condition (B). For this, we shall use the following Runge theorem for conformal minimal immersions (see [3, Theorem 5.3] or [5, Theorem 3.6.1]), together with further results in the recently developed theory of approximation for minimal surfaces.
@@
Let us now proceed to prove items (i) to (ix) in Theorem 1.2. This shall complete the proof of the theorem.
@@
where $\mathrm{dist}_{u}^{\mathfrak{g}}$ denotes the distance on $M$ associated to the pull-back metric $u^{*}\mathfrak{g}$. Let $C=bK_{j}\in E$, $u\in{\rm CMI}(M,\mathbb{R}^{n})$, and $\mu>0$. The Cauchy estimates (relying on the fact that the uniform convergence on a compact set of harmonic or holomorphic functions implies the one of all their derivatives) guarantee the existence of a number $\sigma>0$ with the following property: for any $v\in{\rm CMI}(M,\mathbb{R}^{n})$ with $\|v-u\|_{K_{j}}<\sigma$ the metric $v^{*}\mathfrak{g}$ is so close to $u^{*}\mathfrak{g}$ on $K_{j}$ that $|\mathrm{dist}_{v}^{\mathfrak{g}}(p_{0},C)-\mathrm{dist}_{u}^{\mathfrak{g}}(p_{0},C)|<\mu$. This proves Claim 2.3 (A).
Let $i\in\mathbb{N}$ and fix $u$, $K$, and $\epsilon$ as in Claim 2.3 (B). Choose $C=bK_{j}\in E$, where $j\in\mathbb{N}$ is so large that $K\subset\mathring{K}_{j}$, and fix a number $\delta>0$ to be specified later. By [2, Lemma 4.1] (see also [5, Lemma 7.3.1]) there is a conformal minimal immersion $\hat{u}:K_{j}\to\mathbb{R}^{n}$ such that $\|\hat{u}-u\|_{K_{j}}<\epsilon/2$ and $\mathrm{dist}_{\hat{u}}(p_{0},C)>\delta$, where $\mathrm{dist}_{\hat{u}}$ denotes the distance function on $M$ associated to the metric $\hat{u}^{*}{ds^{2}}$ induced on $M$ via $\hat{u}$ by the standard Euclidean metric $ds^{2}$ in $\mathbb{R}^{n}$. Since $ds^{2}$ and the given metric $\mathfrak{g}$ are comparable in the compact set $\{x\in\mathbb{R}^{n}\colon\mathrm{dist}(x,u(K_{j}))\leq\epsilon\}\subset\mathbb{R}^{n}$, we can choose $\delta>0$ so large that
@@
$\mathcal{F}_{l}(u,C)=\left\{\begin{array}[]{ll}\displaystyle\frac{1}{\mathrm{dist}(u(C),Z)}&\text{if }l=3\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\sup_{C}|\mathcal{K}_{u}|(\rho\circ u)&\text{if }l=4\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\int_{C}|\mathcal{K}_{u}|(\rho\circ u)dA_{u}&\text{if }l=5\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\int_{C}(\rho\circ u)dA_{u}&\text{if }l=6\;.\end{array}\right$.
@@
$\int_{D}|\mathcal{K}(\tilde{u}^{5})|dA_{\tilde{u}^{5}}>i/\delta,\quad{\rm Area}(\tilde{u}^{6}(D))=\int_{D}dA_{\tilde{u}^{6}}>i/\delta.$
@@
Let $M$ be an open Riemann surface and $B\subset\mathbb{R}^{3}$ be an open ball. Reasoning as in the proof of Corollary 1.3 (iv), it suffices to show that the set $X$ of all immersions $u\in{\rm CMI}(M,\mathbb{R}^{3})$ with $u^{-1}(B)\neq\varnothing$ and having infinite index of stability on $u^{-1}(B)$ is a dense $G_{\delta}$ subset. For this, take an exhaustion $K_{j}$, $j\in\mathbb{N}$, of $M$ as in (2.1), and for each $i\in\mathbb{N}$ set
@@
(D) Let $R$ be a compact bordered Riemann surface, and assume that $R$ appears recurrently in $M$, meaning that for any Runge compact set $K\subset M$ there is a smoothly bounded compact domain $L\subset M\setminus K$ being biholomorphic to $R$ such that $K\cup L$ is Runge. Arguing as in the proofs of Theorem 1.2(ii) and Corollary 1.3(ii), one can see that the subspace $\mathrm{CMI}_{R}(M,\mathbb{R}^{n})\subset\mathrm{CMI}(M,\mathbb{R}^{n})$ of immersions rebuilding every conformal minimal immersion $R\to\mathbb{R}^{n}$ is residual (cf. Definition 1.1(c)). Likewise, if $R_{j}$, $j\in\mathbb{N}$, is a sequence of compact bordered Riemann surfaces appearing recurrently in MM, [^4] By conformal surgery theory, one can construct open Riemann surfaces $M$ such that $R_{j}$ appears recurrently in $M$ for all $j\in\mathbb{N}$. then $\bigcap_{j\in\mathbb{N}}\mathrm{CMI}_{R_{j}}(M,\mathbb{R}^{n})$ is still residual in $\mathrm{CMI}(M,\mathbb{R}^{n})$.
@@ Acknowledgements
This research was partially supported by the State Research Agency (AEI) via the grants no. PID2020-117868GB-I00 and PID2023-150727NB-I00, and the “Maria de Maeztu” Excellence Unit IMAG, reference CEX2020-001105-M, funded by MICIU/AEI/10.13039/501100011033 and ERDF/EU, Spain.